OpenStudy (anonymous):

How would i solve for the sec and cot? if given sec(-pi/6)

OpenStudy (anonymous):

you want \(\sec(-\frac{\pi}{6})\) find \(\cos(-\frac{\pi}{6})\) and take the reciprocal

OpenStudy (anonymous):

again find \(-\frac{\pi}{6}\) on the unit circle. first coordinate is cosine, reciprocal is secant

OpenStudy (anonymous):

so secant,cotangent, and co secant are just reciprocals of sin cos and tangent?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but be careful secant is the reciprocal of COSINE and cosecant is the reciprocal of sine

OpenStudy (anonymous):

how would i solve for this equation ? 2xsquared-6x-9=0? i tried using the quadratic formula but im coming up with a wrong answer.

OpenStudy (anonymous):

\[2x^2-6x-9=0\] like that ?

OpenStudy (anonymous):

\[x=\frac{6\pm\sqrt{6^2-4\times 2\times (-9)}}{2\times 2}\] whatever that is

OpenStudy (anonymous):

yes thanks, before i leave i have ACT's tmmrw any helpful tips?